### v1

# BEAM: Bridging Physically-based Rendering and Gaussian Modeling for Relightable Volumetric Video

Abstract
Volumetric video enables immersive experiences by capturing dynamic 3D scenes, enabling diverse applications for virtual reality, education, and telepresence. However, traditional methods struggle with fixed lighting conditions, while neural approaches face trade-offs in efficiency, quality, or adaptability for relightable scenarios. To address these limitations, we present BEAM, a novel pipeline that bridges 4D Gaussian representations with physically-based rendering (PBR) to produce high-quality, relightable volumetric videos from multi-view RGB footage. BEAM recovers detailed geometry and PBR properties via a series of available Gaussian-based techniques. It first combines Gaussian-based performance tracking with geometry-aware rasterization in a coarse-to-fine optimization framework to recover spatially and temporally consistent geometries. We further enhance Gaussian attributes by incorporating PBR properties step by step. We generate roughness via a multi-view-conditioned diffusion model, and then derive AO and base color using a 2D-to-3D strategy, incorporating a tailored Gaussian-based ray tracer for efficient visibility computation. Once recovered, these dynamic, relightable assets integrate seamlessly into traditional CG pipelines, supporting real-time rendering with deferred shading and offline rendering with ray tracing. By offering realistic, lifelike visualizations under diverse lighting conditions, BEAM opens new possibilities for interactive entertainment, storytelling, and creative visualization.
Figure 1: We present BEAM, a novel pipeline that bridges 4D Gaussians with accurate physically-based rendering to produce relightable volumetric videos, delivering immersive and realistic experiences on platforms such as VR.
[…]
1 Introduction
Volumetric video captures dynamic 3D scenes from multiple angles, allowing interactive viewing from any perspective. This technology is crucial for creating immersive experiences in virtual and augmented reality, enhancing storytelling, education, cultural preservation, and telepresence with lifelike, interactive content. However, traditional volumetric video is often limited by fixed lighting conditions captured during recording, which can clash with dynamic or virtual environments, reducing realism and flexibility. Relightable volumetric video overcomes this limitation by enabling post-capture relighting. This allows for seamless integration into dynamic lighting environments and offers creative control over visual aesthetics.
The prevailing workflow [7, 14, 36, 19] for producing relightable volumetric videos in the industry still relies on tracked mesh sequences and texture videos, which can be seamlessly integrated into standard CG pipelines to support relighting under various lighting conditions. However, the intricate reconstruction process often introduces artifacts such as holes and noise, and the quality of relighting remains constrained, frequently resulting in visible imperfections. Neural advancements [56, 81, 74] focus on enabling relighting capabilities using neural factorization within implicit MLPs representations. However, these approaches often face challenges in balancing training efficiency, rendering speed, and output quality, ultimately failing to deliver satisfactory results. Recently, 3D Gaussian Splatting (3DGS)[35], an efficient point-based representation, has achieved photo-realistic rendering at unprecedented frame rates. While dynamic variants[30, 38, 65] can produce high-quality volumetric videos, they fail to produce the detailed geometry necessary for essential operations like relighting. Although efforts [18, 54, 40, 32] have been made to integrate physically-based rendering into the 3DGS pipeline, these methods are often computationally expensive and limited to static scenarios. These limitations severely restrict their applicability in industrial workflows, hindering the efficient production of 4D content.
In this paper, we introduce BEAM, a novel pipeline that bridges 4D Gaussians with accurate physically-based rendering (PBR) for producing relightable volumetric videos from multi-view RGB footage. Our key idea is to robustly recover detailed geometry and decouple the PBR properties (e.g., ambient occlusion, roughness, and base color) using a carefully selected suite of techniques, i.e., rasterization, performance tracking, and ray tracing, all within a Gaussian-based paradigm. As a result, BEAM enables lifelike dynamic scenes that can be seamlessly and CG-friendly integrated into various platforms under diverse lightings (see Fig. 1).
We first recover detailed and spatial-temporally consistent geometries from multiview video input, which organically combines the Gaussian-based performance tracking [30] with the geometry-aware Gaussian rasterization [78]. While the former excels at motion tracking and the latter at static geometry recovery, we unify them in a coarse-to-fine optimization framework. Specifically, we employ coarse joint Gaussians to track non-rigid motion and dense skin Gaussians to preserve intricate geometry details. We adopt a robust optimization process that integrates normal consistency, photometric consistency, and temporal regularization to enhance geometric accuracy and smoothness. This enables accurate depth and normal recovery from the dense Gaussians using the geometry-aware rasterizer [78], providing a robust foundation for material decomposition and relighting.
We further decouple the dense 4D Gaussians to recover detailed material properties, enabling high-quality physically-based rendering grounded in the rendering equation [33] and simplified Disney BRDF [2]. Assuming human-centric scenes with negligible metallic components, we focus on accurately associating roughness, ambient occlusion (AO), and base color properties with the Gaussians, ensuring realistic and adaptable rendering under diverse lighting conditions. To achieve this, we adopt a step-by-step approach to disentangle these properties. Specifically, we first generate a roughness texture using the material diffusion module in previous work [79] with multi-view conditioning, which is associated with the dense Gaussians through UV projection. Then, for the AO and base color, we adopt a 2D-to-3D strategy, where these attributes are estimated in the input views to bake 2D material maps, and then optimize into the corresponding dense Gaussians in the 3D space. This strategy effectively reduces noise and smooths the disentanglement to improve relighting quality. For further 2D AO and based color decomposition, the lighting environment during capturing can be estimated using off-the-shelf tool [47], while the geometry attributes and roughness are obtained in previous stages. Thus, by carefully re-examining and simplifying the rendering equation [33], we identify a critical insight: both 2D AO and base color can be accurately derived by accumulated visibility information for specific points along specific directions during ray tracing. We tailor the Gaussian-based ray tracer [45] to compute such visibility, with a novel alpha blending strategy based on the dense Gaussians. This strategy efficiently captures visibility information, forming the foundation for estimating AO and base color maps in the input viewpoints.
Once the material properties are baked into our dense dynamic Gaussians, these 4D assets seamlessly integrate with traditional CG engines, supporting flexible rendering workflows. For real-time rendering, we adopt deferred shading to deliver immersive and efficient visualizations, while offline rendering leverages ray tracing to precisely capture complex shadows and occlusions. We further develop a Unity plugin enabling seamless integration of 4D assets into various platforms for real-time, lifelike interactions under diverse lighting conditions. This innovation opens new possibilities for storytelling, interactive entertainment, and creative visualization, offering users an immersive journey into dynamic, relightable volumetric worlds.
Figure 2: We propose a novel BEAM method to produce Gaussian sequences. We first use tracking results of joint Gaussians and a normal regularizer to obtain our 4D Gaussians with consistent geometry. Then we infer roughness using a generative model and apply ray tracing to compute the 2D base color and AO maps, which are then used to optimize the corresponding Gaussian attributes. Our results can be rendered under varying lighting conditions using both real-time and offline rendering.
2 Related Work
Human Modeling
In the field of Human Modeling, numerous methods [57, 60, 82, 17] have been proposed to address these challenges. Li [37] integrates temporal denoising into non-rigid mesh template tracking to capture detailed geometry, while High-quality FVV [7] compactly represents human performance using tracked mesh sequences and textured video.
With the advancement of neural rendering techniques [44], several works have incorporated time as a latent variable to handle dynamic scenes [50, 70, 62]. Meanwhile, other approaches [49, 66, 84, 22, 58, 85, 48] leverage the human parametric model [41] to reconstruct the animatable avatar. Furthermore, some hybrid methods [75, 27, 28] combine explicit volumetric fusion with implicit neural techniques to capture more details. Additionally, several methods [64, 26, 28, 53, 55] significantly accelerate training and rendering speed, by leveraging advanced data structures such as voxel grids [16], hash tables [46], and tensor decomposition [4]. Recently, 3DGS [35] ensures both high quality and fast rendering, while dynamic variants [59, 69, 42, 24, 73, 31, 15, 72, 38] enable complex 4D scene reconstruction for advanced human modeling. DualGS [30] uses joint and skin Gaussians to capture motion and detailed appearance. However, these approaches fail to recover detailed geometry and do not support relighting.
Human Relighting
Human Relighting aims to manipulate the reflectance field of the human surface, enabling an immersive fusion with novel illumination. Conventional methods [3, 9, 10, 11, 21, 23, 67, 68] use LightStage systems to capture human reflectance characteristics, requiring costly controlled lighting and dense camera arrays that are not widely accessible. In 2D image-based relighting tasks, previous methods [34, 61]rely on convolutional networks for inference, while recent diffusion-based approaches [13, 77] leverage generative models to learn complex light interactions across large datasets. However, the lack of a 3D representation makes it challenging to maintain 3D consistency of lighting. In addition, some neural relighting methods [1, 81, 76] use NeRF [44] for 3D reconstruction of material properties and lighting effects. However, their rendering quality is inadequate and hard to integrate with traditional CG pipelines. Recent methods [18, 32, 40, 20] leverages 3DGS representation for relighting due to its ability to reconstruct fine details and interaction in CG engines. For human performance relighting, researchers [6, 71, 39, 5, 83, 43] extend mesh-based and neural relighting methods by incorporating body pose priors [41, 29]. However, avatars relying on skeletal priors struggle with complex clothing, wrinkles, and human-object interactions.
3 Method
Given multi-view video inputs, we aim to produce relightable 4D Gaussian sequences with physically-based rendering (PBR) materials, enabling realistic rendering under various lighting. PBR typically requires multiple material components, such as base color, metallic, roughness, normal, and ambient occlusion (AO), as shown in Fig.3. In our approach, as the metallic attribute is negligible for human bodies, we set it to zero and focus on optimizing the remaining properties: normal, roughness, AO, and base color. The complete pipeline is illustrated in Fig.2.
3.1 Gaussian Modeling and Geometry Optimization
To obtain temporally consistent geometric information (depth and normal maps) for physical-based rendering, we seamlessly combine the Gaussian-based performance tracking [30] with the geometry-aware rasterizer [78] within a coarse-to-fine optimization framework. Specifically, our framework uses a dual Gaussian representation to separately model global motion and visual appearance through joint and skin Gaussians. Each skin Gaussian is anchored to multiple joint Gaussians and is warped across frames based on the tracking results of these joint Gaussians. The optimization process integrates a photometric loss $E_{\text{color}}$, a rigid term $E_{\text{rigid}}$, and a temporal regularization term $E_{\text{temp}}$.
To further enhance the optimization of Gaussian geometry, we introduce an additional normal consistency loss $E_{\text{normal}}$. During rasterization, we assume that the intersection points between rays and Gaussians correspond to the maxima in Gaussian values. The depth of a Gaussian is defined as the depth of the intersection point, while the normal of the intersection plane is taken as the Gaussian normal. This rasterizing process generates precise depth maps and normal maps $N_{r}$. Using the obtained depth map, we compute the normal map $N_{d}$ based on a local plane assumption and then measure the normal consistency loss $E_{\text{normal}}$ between $N_{d}$ and $N_{r}$ as follows:
$E_{\text{normal}}=\sum_{i}\omega_{i}(1-N_{r}^{\top}N_{d}),$ (1)
where $i$ indexes the intersected splats along the ray, and $\omega_{i}=\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ represents the blending weight of the intersection point.
The total energy term in our 4D Gaussian modeling and geometric optimization framework is expressed as:
$E=\lambda_{\text{color}}E_{\text{color}}+\lambda_{\text{smooth}}E_{\text{smooth}}+\lambda_{\text{temp}}E_{\text{temp}}+\lambda_{\text{normal}}E_{\text{normal}},$ (2)
During dynamic training, $E_{\text{norm}}$ is introduced after the appearance optimization over 7,000 iterations, followed by an additional 5,000 iterations dedicated to optimizing the normals. For other energy terms not discussed here, please refer to DualGS [30]. The hyperparameters are set as follows: $\lambda_{\text{color}}=1$, $\lambda_{\text{smooth}}=0.001$, $\lambda_{\text{temp}}=0.00005$, $\lambda_{\text{normal}}=0.03$.
Figure 3: Illustration of Gaussian properties. We showcase the normal, roughness, AO and base color maps required for relighting, which are rasterized from our relightable 4D Gaussians.
3.2 PBR Materials Decomposition
To enable physical-based rendering within the Gaussian-based paradigm, we extend 3D Gaussian attributes by incorporating a roughness value and two additional third-order spherical harmonic (SH) properties for AO and base color. And we utilize a step-by-step approach to disentangle these properties.
Roughness.
To efficiently assign a roughness value to each Gaussian, we utilize a generative model that offers significantly faster performance compared to inverse rendering methods. Specifically, we feed the canonical mesh and multi-view images into the Material Diffusion module of CLAY [79] to generate a roughness texture for the mesh. Each valid pixel of the texture is then mapped back to the world coordinate system, allowing us to assign each Gaussian the roughness value of its nearest pixel.
Figure 4: We present the results of different rendering techniques. Real-time rendering focuses on efficiency, while offline rendering delivers more realistic shadows and occlusion effects.
Figure 5: Gallery of our results. We present some real-time rendering results under HDRI settings, which deliver high-fidelity rendering of human performances across challenging motions and complex clothing textures.
Ambient Occlusion and Base Color
For AO and base color, we first estimate these two attributes in the input views for each frame to bake 2D material maps, and then optimize into the corresponding dense skin Gausssians.
Ambient Occlusion $\mathcal{A}(x)$ is an approximation of global illumination, which models the diffuse shadows produced by close, potentially small occluders within a constrained computational budget:
$\mathcal{A}(x)=\frac{1}{\pi}\int_{\Omega}V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (3)
where $V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)$ is the visibility term at 3D point $x$ in direction $\boldsymbol{\omega}_{i}$, $\mathbf{n}$ is the normal of the surface at point $x$, $\Omega$ is the hemisphere centered in $x$ and having $\mathbf{n}$ as its axis.
For the base color, we use a simplified Disney BRDF model [2] composed of a Lambertian diffuse term and a Cook-Torrance specular term [8], with the outgoing radiance $L_{o}$ being a linear combination of these two components. Since for dielectric materials the diffuse part is proportional to the base color while the specular part is independent of it, the rendering equation [33] can be written as:
$L_{o}\left(x,\boldsymbol{\omega}_{o}\right)=\rho(x)L_{o}^{D}\left(x,\boldsymbol{\omega}_{o}\right)+L_{o}^{S}\left(x,\boldsymbol{\omega}_{o}\right),$ (4)
where $L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ is the outgoing radiance at $x$ in direction $\boldsymbol{\omega}_{o}$, computed by mapping image pixel colors to linear space. $\rho$ is the base color, $L_{o}^{D}$ is the diffuse part residue and $L_{o}^{S}$ is the specular part. To simplify the rendering equation, our computation of the base color disregards the indirect illumination effects caused by surface reflections on the human body. Consequently, $L_{o}^{D}$ and $L_{o}^{S}$ are expressed as follows:
$L_{o}^{D}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\frac{1}{\pi}\int_{\Omega}\left(1-F\right)\mathcal{L}(x,\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (5)
$L_{o}^{S}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\int_{\Omega}f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)\mathcal{L}(x,\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (6)
where $\mathcal{L}(x,\boldsymbol{\omega}_{i})=V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})$, $F$ is the approximated Fresnel term, $f_{r_{s}}$ is the specular term in the BRDF, $L_{i}^{\text{env}}$ can be queried from the environment map. To capture the environment map of our multi-view, well-lit dome setup, we position a DSLR camera at the center and take bracketed exposure photographs from multiple directions. These photographs are then processed using PTGui [47] to generate a high dynamic range (HDR) panoramic image, enabling precise calculation of the incoming radiance $L_{i}^{\text{env}}$ from the environment.
Figure 6: Qualitative comparisons of our method against R-3DGS [18], GS-IR [40] and D-2DGS [80]. D-2DGS refers to the results of inverse rendering applied to mesh sequences obtained from D-2DGS to decouple PBR materials. Our method achieves the highest relighting quality. For more detailed comparison results, please refer to the appendix.
We observe that both 2D AO and base color need the visibility term $V^{\text{env}}$. To accurately compute $V^{\text{env}}$, we adopt Gaussian ray tracer from 3DGRT [45], and compute intersections based on the maximum Gaussian response. For efficient computation, a proxy icosahedron mesh is employed to leverage hardware acceleration, with a two-level BVH constructed at both the mesh and instance levels. We accumulate the alpha values of the Gaussians at the intersection points without sorting them:
$\alpha=\sum_{i=1}^{N}O_{i}G_{i}(x_{i}),$ (7)
where $O_{i}$ is the opacity of the $i_{th}$ Gaussian, $G_{i}(x_{i})$ is the response of the Gaussian kernel at $x_{i}$ which makes the response be the maximum value along the ray. We set $V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)=0$ only when $\alpha$ is greater than a threshold $T$ (we choose $T$ as 0.9999); otherwise, $V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)=1$. Ray tracing stops once the accumulated alpha value exceeds the threshold $T$. To avoid occlusion from Gaussians directly above the original depth, we offset the ray origin $o$ along the normal direction: $o=x+\epsilon\mathbf{n}$.
Finally, we use the Monte Carlo method to solve the integrals in Eq. 356. We render the AO image with 50 samples per pixel and the base color image with 100 samples per pixel, then apply Intel Open Image Denoise [25] for denoising.
To optimize the AO and base color material maps into the corresponding dense skin Gaussians, we employ the generated training view material maps as ground truth supervision, while keeping all other geometry-related attributes fixed. The optimization process commences by initializing the AO attributes with zeros and the base color attributes with RGB values. We then train 5000 iterations to obtain both AO and base color attributes separately. Besides, we incorporate the same regularization term to the base color to ensure temporal consistency.
3.3 Physically Based Rendering
By leveraging our relightable 4D Gaussians, we can seamlessly integrate the 4D assets into traditional CG engines, supporting both real-time and offline rendering workflows. For real-time rendering, we utilize the deferred shading to deliver immersive and efficient visualization across diverse settings. For offline rendering, we employ ray tracing, which excels in handling shadows and occlusion relationships, ensuring high-quality results.
Real-time Rendering.
We implement real-time rendering using deferred shading techniques [12], based on the High Definition Render Pipeline (HDRP) in Unity. Specifically, we rasterize our 4D Gaussian sequence with extra PBR attributes, including base color, AO, normals, roughness, and depth maps, and store them in the GBuffer. This GBuffer is then integrated into the original forward transparent stage of HDRP for rendering semi-transparent objects. We also leverage shadow mapping from the HDRP rendering pipeline to perform shadow calculations for Gaussians under different light types. Our approach enables real-time rendering at 100 FPS in 1080P for volumetric videos.
Offline Rendering.
For offline rendering, we sort the intersection points of the Gaussian during ray tracing, and use alpha blending to determine the exact intersection with the entire Gaussian object, simultaneously acquiring the normal and material attributes at that point. This allows full compatibility with the widely used path tracing pipeline. To demonstrate the relighting quality of real-time rendering and offline rendering techniques, we showcase results for both under an environment map, as depicted in Fig. 4. Notably, the offline rendering effectively handles the shadows caused by occlusions in the Gaussian representations.
4 Experiments
To demonstrate our relighting capabilities, we capture human performances with rich textures and challenge body movements using 81 calibrated industrial cameras, enabling high-quality rendering under diverse lighting conditions. Our pipeline is implemented based on 3DGS [35] and trained on a single NVIDIA GeForce RTX 3090. Our method achieves a processing time of 12 minutes per frame for 4D Gaussians modeling and optimization and 4 minutes for materials estimation and baking. The generated relightable 4D Gaussian sequences are fully compatible with VR platforms and CG engines, enhancing immersive experiences during playback and editing.
4.1 Comparison
We compare our method against state-of-the-art static Gaussian relighting techniques, R-3DGS [18], GS-IR [40], both of which reconstruct geometry frame by frame. For dynamic comparison, we first obtain the mesh sequences using D-2DGS [80] and then apply the path replay backpropagation method [63] for inverse rendering to decouple attributes, such as base color and roughness for each frame individually. As shown in Fig. 6, the normals and AO decoupled by Relightable-3DGS are blurry, resulting in significant relighting artifacts. GS-IR struggles to reconstruct smooth normals and effectively separate the base color from other attributes. While D-2DGS with inverse rendering produces smoother normals and more accurate attribute decoupling, the relighting results are prone to losing high-frequency details. In contrast, our method produces smooth normals and accurately decouples the AO and base color, enabling high-fidelity relighting results.
For quantitative comparison, we conduct evaluations on synthetic data to generate ground truth images under predefined lighting conditions. We use Blender to simulate the similar capture perspectives of our dome system and render human meshes from RenderPeople[51] into corresponding viewpoints using the CYCLES engine. we employ PSNR, SSIM, and LPIPS as metrics. To ensure a fair and precise comparison, we compute the metrics for AO, base color, and relighting results across two synthetic sequences, each consisting of 150 frames. In addition, we compute these metrics within the bounding box of the human region. As shown in Tab. 1, our method surpasses the other techniques in all metrics evaluated.
Table 1: Quantitative comparison with SOTA relighting methods on our blender-rendered synthetic dataset. Green and yellow cell colors indicate the best and the second-best results.
AO Base Color Relighting
[…]
R-3DGS 20.67 0.791 0.307 20.65 0.853 0.166 24.66 0.855 0.105
GS-IR 18.95 0.850 0.345 13.35 0.702 0.252 23.05 0.858 0.177
D-2DGS 24.19 0.906 0.232 21.04 0.839 0.145 23.23 0.857 0.198
Ours 25.32 0.924 0.168 21.47 0.906 0.084 25.57 0.895 0.086
Figure 7: Ablation on parameter settings for materials decomposition. Our full model provides accurate AO and base color representations, resulting in high-quality relighting with minimal artifacts.
Figure 8: Ablation study on the number of geometry-aware Gaussians. With $\sim 140,000$ Gaussians, BEAM achieves high rendering quality and geometric reconstruction accuracy, while ensuring efficient training and rendering.
4.2 Evaluation
Materials Decomposition.
We conduct a qualitative ablation study on materials decomposition to the impact of different variables on AO, base color, and the final relighting results. For the offset variable in the Gaussians surface estimation during ray tracing, we present the results without the offset and with an excessively large offset (0.1), in the first and second columns of Fig. 7. We found that without an offset, AO and base color contained many black artifacts, and the Gaussian surface estimation suffered significant degradation. On the other hand, an offset of 0.1 led to overestimation, causing incorrect lighting decoupling and producing overly bright artifacts. Additionally, regarding the SH order required to represent AO and base color, as illustrated in the third column of Fig. 7, using a zero-order SH Gaussian resulted in blurriness. In contrast, our full ray tracing and material optimization pipeline, with an offset set to 0.02 and using third-order SH degree with geometry-aware Gaussians, provides accurate calculations and representations of the Gaussian’s AO and base color attributes.
The Number of Gaussians.
We evaluate the impact of varying the number of geometry-aware Gaussians on both rendering quality and mesh quality across 100 frames in our virtual dataset, using PSNR and Chamfer Distance respectively. As shown in Fig. 8, using $\sim 140,000$ geometry-aware Gaussians strikes an optimal balance between rendering quality and geometric reconstruction accuracy, without introducing excessive redundancy. This point count ensures efficient training and is well-suited for applications such as VR and AR.
4.3 User Study
We conduct a user study to evaluate the temporal reconstruction quality and normal consistency of our 4D Gaussians. We assess our method on our captured dataset, comparing it against R-3DGS, GS-IR, as well as D-2DGS, and ask 30 users to select the best option. For GS-IR and R-3DGS, we train Gaussian models for 200 frames individually and concatenate the per-frame normal maps. For our method and D-2DGS, we rasterize a normal map sequence from a 200-frame dynamic Gaussian sequence. In terms of temporal reconstruction quality, 95.65% of users prefer our method, while 87% choose our approach for normal consistency. These preference results clearly indicate a significant advantage of our method over the competing approaches, demonstrating its superior performance in both aspects.
5 Conclusion
Limitations.
Although our method achieves high-quality immersive rendering, there are some limitations. First, the integration of new attributes into the original Gaussian expression will lead to substantial storage requirements for dynamic Gaussian sequences. Future work will consider compressing these sequences. Second, we approximate the rendering equation to obtain the 2D material maps, which introduces errors in our decoupled material results and does not accurately reflect the real physical world. Future work may address this issue by incorporating large models like video generation. Furthermore, our approach, focusing on relighting for a dynamic reconstruction sequence, does not support pose-driven animation or the generation of new poses. Future work will focus on optimizing these aspects to improve robustness and applicability.
We have presented a Gaussian-based approach for reconstructing detailed geometry and PBR materials to produce relightable volumetric videos. We employ a coarse-to-fine training strategy and effective geometric constraints to accurately model the dynamic geometry of 4D Gaussians. Additionally, we decouple PBR materials by using ray tracing to compute the lighting effects and obtain base color and AO maps, while leveraging generative methods to infer roughness. These materials are then baked into the corresponding attributes of the Gaussians. With deferred shading and ray tracing techniques, our Gaussian sequence supports both efficient real-time rendering and more realistic offline rendering. Experimental results demonstrate the advantages of our approach in generating high-quality dynamic normal maps and material decomposition, as well as its relightability under a variety of lighting conditions. Our method is highly compatible with traditional CG engines, offering significant potential for enhancing rendering realism and flexibility.
Appendix A Energy Term
To model human motion using Gaussian sequences, we first employ the color energy $E_{\text{color}}$ during the training process:
$E_{\text{color}}=(1-\lambda_{\text{color}})\mathcal{L}_{1}+\lambda_{\text{color}}\mathcal{L}_{\text{D-SSIM}},$ (8)
where $\mathcal{L}_{1}$ is the $L_{1}$ photometric loss and $\mathcal{L}_{\text{D-SSIM}}$ is the D-SSIM term. We then employ a smooth regularizer to constrain the joint Gaussians motion locally as-rigidas-possible(ARAP):
$E_{\text{smooth}}=$ $\sum_{i}\sum_{k\in\mathcal{N}(i)}w_{i,k}\left\|R\left(q_{i,t}^{j}*{q_{i,t-1}^{j}}^{-1}\right)\right.$ (9)
$\quad\left.\left(p_{k,t-1}^{j}-p_{i,t-1}^{j}\right)-\left(p_{k,t}^{j}-p_{i,t}^{j}\right)\right\|_{2}^{2},$
where $\mathcal{N}(i)$represents the set of neighboring joint Gaussian kernels of $i$, $R(\cdot)$ converts quaternion back into a rotation matrix and $w_{i,k}$ corresponds to the blending weights according to DualGS [30]. Furthermore, we apply the temporal regularizer $E_{\text{temp}}$ to promote coherent appearances by constraining Gaussian attributes ($C_{i,t},\sigma_{i,t},s_{i,t}$) to be consistent with the previous frame:
$E_{\text{temp}}=\sum_{a\in\{\mathcal{C},\sigma,s\}}\lambda_{a}\|a_{i,t}-a_{i,t-1}\|_{2}^{2},$ (10)
Together with $E_{\text{normal}}$ we mentioned above, the overall energy term in our 4D Gaussian modeling is expressed as:
$E=\lambda_{\text{color}}E_{\text{color}}+\lambda_{\text{smooth}}E_{\text{smooth}}+\lambda_{\text{temp}}E_{\text{temp}}+\lambda_{\text{normal}}E_{\text{normal}},$ (11)
which helps us obtain accurate geometric information while modeling human performance sequences.
Appendix B Base Color Decomposition
To compute base color, we ignore the effect of indirect illumination caused by reflections on the surface of the human body, the rendering equation can be approximated as:
$L_{o}\left(x,\boldsymbol{\omega}_{o}\right)\approx\int_{\Omega}f_{r}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (12)
where $x$ and $\mathbf{n}$ are the surface point and its normal vector, $f_{r}$ is the Bidirectional Reflectance Distribution Function (BRDF), $L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ denotes the outgoing radiance in direction $\boldsymbol{\omega}_{o}$.
We decompose the BRDF term $f_{r}$ into two components, the diffuse term $f_{r_{d}}$ and the specular term $f_{r_{s}}$, where we use the Lambertian model for $f_{r_{d}}$ and Cook-Torrance microfacet specular shading model [8] for $f_{r_{s}}$:
$f_{r}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)$ $=f_{r_{d}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)+f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right),$ (13)
$f_{r_{d}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)$ $=\frac{1-F}{\pi}\rho(x),$ (14)
$f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)$ $=\frac{DGF}{4(\mathbf{n}\cdot\boldsymbol{\omega}_{i})(\mathbf{n}\cdot\boldsymbol{\omega}_{o})},$ (15)
where $D$, $G$, $F$ are respectively the GGX normal distribution function (NDF), the geometric attenuation function, and the approximated Fresnel term. We then split $L_{o}$ into two parts:
$L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\int_{\Omega}\left(f_{r_{d}}+f_{r_{s}}\right)\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i},$ (16)
$=\int_{\Omega}f_{r_{d}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}+\int_{\Omega}f_{r_{s}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i},$ (17)
$=\frac{\rho(x)}{\pi}\int_{\Omega}\left(1-F\right)\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}+\int_{\Omega}f_{r_{s}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i},$ (18)
where $\mathcal{L}=V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})$.
The $D$ and $G$ in $f_{r_{s}}$ are both irrelevant with the base color $\rho$. For the Fresnel term $F$, we use the Schlick Fresnel approximation [52]:
$F=F_{0}+(1-F_{0})\left(1-\left(\mathbf{h}\cdot\boldsymbol{\omega}_{o}\right)\right)^{5},$ (19)
where $\mathbf{h}=\frac{\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}}{\left\|\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}\right\|}$, $F_{0}$ represents the specular reflectance at normal incidence and is achromatic for dielectrics and chromatic for metals. Since the human body is mostly composed of dielectric materials, we assume $F_{0}=0.04$, so that $F$ is also irrelevant with $\rho$. Consequently, we can solve the base color $\rho$ from Eq. 18:
$\rho(x)=\frac{\pi\left(L_{o}(x,\boldsymbol{\omega}_{o})-\int_{\Omega}f_{r_{s}}\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}\right)}{\int_{\Omega}\left(1-F\right)\mathcal{L}\mathrm{d}\boldsymbol{\omega}_{i}}.$ (20)
Appendix C Additional Evaluations
In this section, we present more details on the qualitative comparison conducted on synthetic human data, including normal, roughness, base color, AO, and relight rendering effects in Fig. 9.
Appendix D Applications
In Fig. 10, we showcase the application of watching an actor dancing in a scene with varying lighting. Viewers can immerse themselves in a realistic environment with dramatic lighting changes, observing the actor’s dance performance seamlessly integrated into the surroundings. We can capture, process, and generate high-fidelity human 4D assets with our dome cameras and BEAM pipeline, reproducing outstanding performances in different scenes and lightings.
At the same time, our compatibility with the traditional CG pipeline allows us to edit scenes in the Unity plugin we developed or integrate multiple human performances, as shown in Fig. 11. We can align multiple performances through CG platform, edit the scene and lighting, and present stunning visual effects of the performances.
Figure 9: We present additional comparison results on synthetic humans, demonstrating the high-quality relighting capabilities of our method.
Figure 10: We demonstrate our VR application on PICO, enabling an immersive experience with relightable volumetric video.
Figure 11: We showcase our custom Unity plugin for relightable 4D Gaussian assets. Our relightable 4D Gaussians are CG-friendly, supporting scene and lighting environment editing within Unity.